OpenStudy (rebeleco518):

Need help understanding 5-10 8th grade math questions.

OpenStudy (kidrah69):

Post ONE question per post what is the question

OpenStudy (rebeleco518):

Yes, ok, here is one...

OpenStudy (rebeleco518):

How many solutions does the equation 4x + 8 = −2 − 2 + 5x have?  -One  -None  -Infinitely many

OpenStudy (kidrah69):

weell combine the like terms

OpenStudy (rebeleco518):

So if I just solve it, I can get the answer? I didn't know if there was an easier way.

OpenStudy (kidrah69):

well solve it then i will show u the trick :)

OpenStudy (rebeleco518):

Okey-dokey, give me a second...

OpenStudy (kidrah69):

Ok so when a solution has the same number equaling the same number it will be infinite solutions example: 0=0 If a solution has 2 different numbers equaling each other it will be no solution such as: 0=1 One solution means ONE SOLUTION like this : X=6

OpenStudy (rebeleco518):

Ok, thank you very much, I got 1 solution for the question, is that correct?

OpenStudy (kidrah69):

yep :) good job!

OpenStudy (rebeleco518):

Ok. All the questions are that way so now that I understand I don't need anymore help! Thank you so very much!!

OpenStudy (kidrah69):

lol im glad i can help! :D

OpenStudy (rebeleco518):

Actually, @kidrah69, I have a question about your reply.

OpenStudy (rebeleco518):

This is the one I have a question about @kidrah69 --> Ok so when a solution has the same number equaling the same number it will be infinite solutions EXAMPLE: 0=0 If a solution has 2 different numbers equaling each other it will be no solution such as: 0=1 One solution means ONE SOLUTION like this : X=6

OpenStudy (rebeleco518):

You say in the first 2 when a certain #=#, that will tell you how many solutions there will be. But in the equations I am doing, x=a certain #. Can you please explain @kidrah69 ?

OpenStudy (kidrah69):

all of them?

OpenStudy (rebeleco518):

just where you say your examples like 0=0 or 0=1. How is there a # instead of a x?

OpenStudy (kidrah69):

Oh ok let me show you an example :)

OpenStudy (rebeleco518):

ok!

OpenStudy (kidrah69):

i hope that clears it up :)

OpenStudy (rebeleco518):

Ooohhhh, Ok! Can you check a few for me?

OpenStudy (rebeleco518):

It's just 4 of this topic @kidrah69

OpenStudy (kidrah69):

sure :)

OpenStudy (rebeleco518):

Ok, thanks! Here's the first one... How many solutions does the equation 5m − 5m − 12 = 14 − 2 have? Zero One -this one? Two

OpenStudy (kidrah69):

\[ 5m − 5m − 12 = 14 − 2 \] \[0m − 12 =12\] add 12 to both sides \[0m=24\] divide \[m=24\] correct :)

OpenStudy (rebeleco518):

Thanks! Here is #2...

OpenStudy (rebeleco518):

Which statement is true about the equation 4.5z = 4.5z − 6.25 + 6.25?    It has one solution.   It has two solutions.    It has infinitely many solutions.  a?

OpenStudy (kidrah69):

can you repost this its lagging for me

OpenStudy (rebeleco518):

ok

OpenStudy (rebeleco518):

Which statement is true about the equation 4.5z = 4.5z − 6.25 + 6.25?   It has one solution.-------this one?   It has two solutions.   It has infinitely many solutions. 

OpenStudy (kidrah69):

\[4.5z = 4.5z − 6.25 + 6.25 \] \[ 4.5z = 4.5z − 6.25 + 6.25~combine~like~terms\] z=8.5 yep thats right

OpenStudy (rebeleco518):

Ok!! here is the last one I believe...

OpenStudy (rebeleco518):

How many solutions does the equation −4y + 4y + 2 = 2 have?  Two  None----this one?  Infinitely many

OpenStudy (kidrah69):

0=0 so which is it? and are these test questions?

OpenStudy (rebeleco518):

It should be infinitely, my mistake. And it was a pre-test, I got these questions wrong and wanted to study and learn the correctly before the actual test. So don't worry @kidrah69 , you're not helping me cheat!!